OpenStudy (anonymous):

What is the sum of all even integers between 99 & 301?

OpenStudy (cwrw238):

sum the arithmetic series with a1 = 100, d = 2, last term = 300

OpenStudy (ajprincess):

100,102..........,300 This is an arithmetic progression. \(T_n=a_1+(n-1)d\) \(300=100+(n-1)2\) \(300-100=2n-2\) \(200=2n-2\) Solve for n. \[S_n=n\frac{ a+l }{ 2 }\] a=100 l=300 Using this find the sum.

OpenStudy (cwrw238):

n = (301 - 99 ) / 2

OpenStudy (anonymous):

my ans is 20200 ,,,is it correct

OpenStudy (ajprincess):

  ya u right:)